OpenStudy (anonymous):

How close does the curve y=2x^2 come to the point (0,3)?

OpenStudy (nincompoop):

very close? not so close, slightly off close

OpenStudy (nincompoop):

solve it :)

OpenStudy (anonymous):

But don't I have to use derivatives to figure out exactly??

OpenStudy (nincompoop):

derivative? sure you can.. but do you really need to?

OpenStudy (anonymous):

I think I do have to use it since that is what we are learning in class

OpenStudy (anonymous):

How would you do this problems using derivatives

OpenStudy (nincompoop):

that's a parabola you are given and the standard-form of a parabola is y = a(x-h)^2 + k given y = 2x^2 means you have zero values for vertex (h,k), which is your lowest point if your vertex is at zero, zero means that you're close three units away from the point (0,3) the rest is up to you to determine whether units up or down :)

OpenStudy (nincompoop):

I didn't do any derivatives nor mathematical calculation. I just based in on the simplest form of a parabola, which is y = x^2 your y = 2x^2 just means that the distance between your x-intercepts is 2 units

OpenStudy (anonymous):

Oh I see I se,what confused me is how to apply derivatives

OpenStudy (nincompoop):

I'd ask and tag the math gods and goddesses of this website. I am just a nincompoop :)